OpenStudy (anonymous):

What is the standard form of an equation for a line that has a slope of -4/5 and passes through the point (2, -10)? 6x + 5y = -38 4x + 5y = -42 4x + 7y = -62 -4x + 5y = -58

OpenStudy (anonymous):

So recall that the general form of a line is\[y=mx+b\]where m is the slope, and b is the y-intercept, alright, now also recall that to find the equation of a line given a point and a slope, you need to use the point-slope-formula which looks like this\[y-y_1=m(x-x_1)\]where m is again, your slope, and the point\[(x_1,y_1)\]is the point that you want the line to go through

OpenStudy (anonymous):

Standard form is just moving all the variables to one side of general formula, and making sure there are no fractions

OpenStudy (amistre64):

and ..... that the first one is the x part and that its not negative

OpenStudy (anonymous):

Oh yeah, dangit, i forgot about that part of the rule, such a simple thing that you get used to i guess

OpenStudy (amistre64):

4 years ago, taking a test ... got it wrong cuase of that simple rule :)

OpenStudy (anonymous):

Ouch, i probably would have made the same mistake, and cried about it

OpenStudy (anonymous):

4x+5y=-42

OpenStudy (anonymous):



OpenStudy (anonymous):

The slope calculation wasn't necessary since the problem gave it to you, but yes, you definitely got the correct answer, great job

OpenStudy (anonymous):

thank you